MsBrains (ms-brains):

1. At a fundraiser breakfast, the bill for three glasses of orange juice and five pancake specials is $7.60, whereas the bill for one glass of orange juice and two pancake specials is $2.90. (a) Write a system of two equations to model this problem. Let j stand for the price of a glass of juice and let p stand for the price of a pancake special. (b) Use the substitution method to solve the system of equations. (c) What would be the bill for one glass of orange juice and one pancake special?

MsBrains (ms-brains):

"three glasses of orange juice and five pancake specials is $7.60" 3j + 5p = 7.60 "one glass of orange juice and two pancake specials is $2.90" j + 2p = 2.90 ::::: Rearrange second equation to solve for j: j = 2.90-2p substitute into first equation: 3(2.90-2p) + 5p = 7.60 8.70 - 6p + 5p = 7.60 p = 1.10 j = 2.90-2p = 0.70 ::::: j + p = 0.70 + 1.10 = 1.80 ^Is that math correct?

OpenStudy (wcrmelissa2001):

yes I think it's correct. Another way to check is substitute your values for orange juice and pancake back into the original one. E.g. 3j+5p = 7.60 3(0.7)+5(1.10)=7.60? If yes, then it's right

MsBrains (ms-brains):

Thank you! Would you mind checking another problem for me?

OpenStudy (wcrmelissa2001):

sure